The letters of the word Mississippi are placed in a hat. What is the probability that the first two letters drawn are both i if the first letter drawn is not replaced?
Answer by stanbon(75887)

The letters of the word Mississippi are placed in a hat. What is the probability that the first two letters drawn are both i if the first letter drawn is not replaced?
----------------------------
# of i's = 4
# of letters = 11
-----
P(i and i) = P(i)*P(i|i) = (4/11)*(3/10) = 12/110 = 6/55
-------------------------------------------------------------
Cheers,
Stan H.